### v1

# Equivariant CM minimization for extremal manifolds

Abstract.
We prove an equivariant version of the CM minimization conjecture for extremal Kähler manifolds. This involves proving that, given an equivariant punctured family of polarized varieties, a relative version of the CM degree is strictly minimized by an extremal filling. This generalizes a result by Hattori for cscK manifolds with discrete automorphism group by allowing automorphisms and extremal metrics. As a main tool, we extend results by Székelyhidi on asymptotic filtration Chow stability of cscK manifolds with discrete automorphism group to the extremal setting.
1. Introduction
A central problem in complex and algebraic geometry is the construction of moduli spaces parameterizing projective varieties. In our case of interest, we consider the moduli problem for polarized varieties; that is, projective varieties endowed with an ample line bundle. The valuative criterion for separatedness ensures that a moduli space of a fixed class of polarized varieties is separated provided that any family $(X,L)\to\Delta^{*}$ of polarized varieties of said class over the open punctured disc $\Delta^{*}$ in $\mathbb{C}$ has at most one filling $(\mathcal{X},\mathcal{L})\to\Delta$ to the whole open disk, up to isomorphism.
This class of problem may be approached numerically: in a geometric invariant theory (GIT) setting, Wang–Xu [WX12] proved that the semistable filling for family over a fixed base minimizes the GIT height, a numerical invariant introduced Wang [Wan12]. It is known to experts that this can be improved to strict minimization for polystable fillings, giving a numerical proof of separatedness of GIT quotients. Wang’s work was motivated by Cornalba–Harris [CH88], Zhang [Zha96] and Xiao [Xia87], and raised the question of how to extend this theory to the setting of stability of varieties.
The analogue of GIT stability for polarized varieties is K-stability [Tia97, Don02], where the numerical invariant generalizing Wang’s GIT height is the degree of the Chow–Mumford (CM) line bundle. In this case, the associated minimization problem is known as the CM minimization conjecture for K-polystable fillings [Wan12, Oda12, WX12]. The CM line bundle is a $\mathbb{Q}$-line bundle defined on the base of flat family of polarized varieties, originally introduced by Fujiki–Schumacher [FS90] in their construction of the moduli space of constant scalar curvature Kähler (cscK) manifolds, and was further developed by Paul–Tian [PT06].
Hattori proved the following result on the degree of the CM line bundle $\mathrm{CM}(\mathcal{X},\mathcal{L})$ for cscK fillings $(\mathcal{X},\mathcal{L})$ over curves:
Theorem 1.1 ([Hat24, Theorem 1.2]).
Let $C$ be a smooth projective curve with a special point 00, and let $(\mathcal{X},\mathcal{L})\to C$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\to C$ be fillings of a flat family $(X,L)\to C\setminus\{0\}$ such that the special fiber $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth, admits a cscK metric and has discrete automorphism group. Then
$\mathrm{CM}(\mathcal{X},\mathcal{L})\leq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}),$
where the inequality is strict if and only if the fillings are not isomorphic.
The existence of cscK metrics is closely related to the stability of polarized manifolds, through the Yau–Tian–Donaldson conjecture [Yau92, Tia97, Don02], which predicts that a polarized manifold admits a cscK metric if and only if it is K-polystable. Thus Theorem 1.1 may be viewed as a proof of the CM minimization conjecture, with an analytic stability hypothesis on the special fiber, under a further discrete automorphism group assumption. The CM minimization conjecture was solved completely algebro-geometrically in the Fano case, by Blum–Xu [BX19, Section 3] and Xu [Xu21, Section 9.4].
As we shall see in our work, one may replace the projective curve with the formal disk as the base, in line with the valuative criterion for separatedness. Since the disk is not proper, the degree of the CM line bundle for a model is not well-defined, but one may identify the difference of their degrees as a rational number. Theorem 1.1 implies separatedness of the moduli space of polarized cscK manifolds with discrete automorphism group, a result originally proven by Fujiki–Schumacher [FS90], and which also follows from Donaldson’s result [Don01] on asymptotic Chow stability of cscK manifolds with discrete automorphism group, using GIT.
1.1. Main results
Let $R$ be a discrete valuation ring (DVR) with fraction field $K$, whose residue field is the field of complex numbers $\mathbb{C}$. The spectrum $\operatorname{Spec}R$ consists of a closed point 00 and a generic point $\operatorname{Spec}K$. The guiding example is the ring of formal power series $R=\mathbb{C}[\![t]\!]$, whose spectrum is called the formal disk and $K=\mathbb{C}(\!(t)\!)$ is the field of formal Laurent series.
Let $(X,L)$ be a polarized $K$-variety, which we think of as a degeneration of a polarized variety over $\operatorname{Spec}K$, and choose a maximal torus $T$ in the automorphism group $\mathrm{Aut}_{K}(X,L)$. The main result of the paper is the following generalization of Theorem 1.1, allowing automorphisms and extremal metrics (i.e. Kähler metrics $\omega$ satisfying ¯$\overline{\partial}\,\mathrm{grad}^{1,0}_{\omega}\mathrm{scal}(\omega)=0$, see [Cal82]):
Theorem 1.2.
Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be $T$-equivariant models of $(X,L)$ such that the restricted torus $T_{0}$ in $\mathrm{Aut}(\mathcal{X}_{0},\mathcal{L}_{0})$ is maximal. If the special fiber $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth and admits an extremal metric, then there exists a rational one-parameter subgroup $\xi$ of $T$ such that
$\mathrm{CM}(\mathcal{X}_{\xi},\mathcal{L}_{\xi})\leq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}),$
where the inequality is strict if and only if for all one-parameter subgroups $\beta$ of $T$, the twisted model $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})$ is not isomorphic to $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$.
The rational one-parameter subgroup $\xi$ may be given explicitly in terms coefficients of Hilbert polynomials (see Remark 4.8). Analytically, extremal metrics arise as critical points of the Calabi functional and as fixed points of the Calabi flow up to biholomorphisms, and are a natural generalization of cscK metrics when automorphisms are involved (see [Szé14, Section 4]). Algebraically, their existence is conjectured to be equivalent to relative K-stability [Szé07], and they are relevant in the understanding of instability stratifications of the stack of polarized varieties.
In the special case of cscK metrics, one obtains the following:
Corollary 1.3.
Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be $T$-equivariant models of $(X,L)$ such that the restricted torus $T_{0}$ in $\mathrm{Aut}(\mathcal{X}_{0},\mathcal{L}_{0})$ is maximal. If the special fiber $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth and admits a cscK metric, then
$\mathrm{CM}(\mathcal{X},\mathcal{L})\leq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}),$
where the inequality is strict if and only if for all one-parameter subgroups $\beta$ of $T$, the twisted model $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})$ is not isomorphic to $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$.
Note that even in the case of discrete automorphisms, Corollary 1.3 is a generalization of Theorem 1.1, as we allow models over DVRs rather than over proper curves. For this, we use a definition of the difference of CM degrees through the Knudsen–Mumford expansion, which we show recovers the definition used by Hattori when the models over $\operatorname{Spec}R$ are the restriction of models over proper curves. Our proof in this generality then involves proving a generalization of the Wang–Odaka intersection-theoretic formula for the Donaldson–Futaki invariant involved in K-stability [Oda13, Wan12] to pairs of models over DVRs (Theorem 2.4).
Theorem 1.2 proves separatedness of the moduli space of extremal manifolds with fixed maximal torus, a result that may also be obtained using Seyyedali’s proof of relative asymptotic Chow stability of extremal manifolds [Sey17] (see also [Mab18, ST21, Has21]); in fact, we will use this in order to prove Theorem 1.2. Both Theorem 1.2 and Corollary 1.3 may be viewed as generalizations of results on relative K-(poly)stability of extremal and cscK manifolds due to Stoppa [Sto09] and Stoppa–Székelyhidi [SS11] to families of polarized varieties.
We also obtain relative arc K-polystability of extremal manifolds:
Corollary 1.4.
Let $(Y,H)$ be an extremal polarized manifold and let $T$ be a maximal complex torus in $\operatorname{Aut}(Y,H)$. Then for all $T$-equivariant arcs $(\mathcal{Y},\mathcal{H})$ for $(Y,H)$, we have
$\mathrm{DF}_{T}(\mathcal{Y},\mathcal{H})\geq 0,$
where the inequality is strict if and only if for all one-parameter subgroups $\beta$ of $T$, the twisted arc $(\mathcal{Y}_{\beta},\mathcal{H}_{\beta})$ is not trivial.
This extends arc K-polystability of cscK manifolds to the extremal setting, proven by Dervan–Reboulet [DR24, Corollary 1.2] (see also [Szé15, Don12] in the absence of automorphisms). The relative Donaldson–Futaki invariant for arcs used in Corollary 1.4 may be defined as the difference of CM degrees between a twist of the arc and the trivial arc, so that Corollary 1.4 follows directly from Theorem 1.2.
1.2. Strategy of the proof
We briefly outline the strategy of proving Theorem 1.2, which is divided into two steps: the first step is proving that, given two equivariant models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, there exists a torus-invariant good filtration $\mathcal{F}$ of $(\mathcal{X}_{0},\mathcal{L}_{0})$, such that
(1.1) $\mathrm{DF}_{T_{0}}(\mathcal{F})=\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})-\mathrm{CM}(\mathcal{X}_{\xi},\mathcal{L}_{\xi}),$
for some rational one-parameter subgroup $\xi$ of $T$. This filtration was constructed by Hattori [Hat24, Theorem 3.5] in the non-equivariant case and generalizes a construction by Blum–Xu [BX19] for degenerations of Fano varieties. The left-hand side of Equation (1.1) is the relative Donaldson–Futaki invariant for invariant good filtrations, generalizing Hattori’s Donaldson–Futaki invariant for good filtrations, and the right-hand side is difference CM degrees of a pair of models over a DVR, both of which we will introduce in the present paper.
The second step is proving a new stability result involving the left-hand side of the Equation (1.1) in the case where $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth and extremal. More precisely, we prove that for all torus-invariant good filtrations $\mathcal{F}$, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$, where the inequality is strict whenever the reduced norm $||\mathcal{F}||_{T_{0}}$ is strictly positive. This generalizes works by Székelyhidi [Szé15], Hattori [Hat24] and Stoppa–Székelyhidi [SS11] for extremal and cscK manifolds with automorphisms.
1.3. Structure of the paper
This paper is divided into three sections. In the second section, we recall the necessary prerequisites on models, CM degrees and filtrations. Notably, we introduce the notion of the difference of CM degrees of a pair of models over a DVR and prove a generalization of the Wang–Odaka intersection-theoretic formula in this setting.
In the third section, we establish the second step of the strategy, proving that an extremal polarized manifold is relatively K-polystable with respect to good filtrations (Theorem 3.2). To prove this, we extend Székelyhidi’s blowup formula for Chow weights of filtrations [Szé15, Page 18] to the relative setting, which involves producing an asymptotic expansion for inner products. We also extend work of Boucksom [Szé15, Section 8] on asymptotic vanishing orders of graded linear series to find a destabilizing invariant point to blow up. Ultimately, we prove that extremal manifolds are relatively asymptotically Chow stable with respect to filtrations.
In the fourth and final section, we prove the first step of the strategy by giving an understanding of the twist of Hattori’s good filtration, and complete the proofs of Theorem 1.2 and Corollary 1.3. Notably, we prove that the twist of the good filtration of two models coincides with the good filtration of the twist of one of the models.
1.4. Acknowledgments
[…]
2. Preliminaries
2.1. The CM line bundle and models
We start with the introduction of the CM line bundle over a general irreducible base, following the treatment in [FR06]. Let $(\mathcal{X},\mathcal{L})\to B$ be a flat family of $n$-dimensional $\mathbb{Q}$-polarized varieties over an irreducible scheme $B$. This consists of a flat morphism $\mathcal{X}\to B$, of relative dimension $n$, and a relatively semi-ample $\mathbb{Q}$-line bundle $\mathcal{L}$ on $\mathcal{X}$. Assume that $n\geq 1$.
By the Knudsen–Mumford expansion [KM76, Theorem 4], there exist $\mathbb{Q}$-line bundles $\lambda_{n+1}(\mathcal{X},\mathcal{L})$ and $\lambda_{n}(\mathcal{X},\mathcal{L})$ on $B$ such that
$\det\pi_{*}(k\mathcal{L})=\frac{k^{n+1}}{(n+1)!}\lambda_{n+1}(\mathcal{X},\mathcal{L})-\frac{k^{n}}{2n!}\lambda_{n}(\mathcal{X},\mathcal{L})+O(k^{n-1})$
as $k\to+\infty$, for sufficiently divisible $k$. Moreover, by flatness, the coefficients of the Hilbert polynomial $h^{0}(\mathcal{X}_{b},k\mathcal{L}_{b})=a_{0}k^{n}+a_{1}k^{n-1}+O(k^{n-2})$ are independent of choice of $b$ in $B$.
Definition 2.1.
The Chow–Mumford (CM) line bundle associated to the family $(\mathcal{X},\mathcal{L})\to B$ is the $\mathbb{Q}$-line bundle
$\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})=\frac{1}{2V(\mathcal{X}_{b},\mathcal{L}_{b})}\bigg{(}\lambda_{n}(\mathcal{X},\mathcal{L})-\frac{n\mu(\mathcal{X}_{b},\mathcal{L}_{b})}{n+1}\lambda_{n+1}(\mathcal{X},\mathcal{L})\bigg{)}$
where $V(\mathcal{X}_{b},\mathcal{L}_{b})=n!a_{0}$ and $\mu(\mathcal{X}_{b},\mathcal{L}_{b})=-\frac{2a_{1}}{na_{0}}$ denote the volume and slope of any fiber $(\mathcal{X}_{b},\mathcal{L}_{b})$.
The definition above is due to Paul–Tian [PT06], while the original construction is due to Fujiki–Schumacher [FS90], where it was used to study families and construct moduli of polarized varieties. In the Fano case, the CM line bundle is known to be ample on the good moduli space of K-polystable Fano varieties, proving its projectivity [XZ20, CP21].
Our definition of the CM line bundle is such that it is invariant under scaling, that is,
$\lambda_{\mathrm{CM}}(\mathcal{X},r\mathcal{L})=\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L}),$
for all integers $r\geq 1$. In the case where the structure morphism $\pi:\mathcal{X}\to B$ admits a relative canonical bundle $K_{\mathcal{X}/B}$, by Grothendiek–Riemann–Roch, the first Chern class of the CM line bundle satisfies
$c_{1}(\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L}))=\frac{\pi_{*}\big{(}c_{1}(\mathcal{L})^{n}c_{1}(K_{\mathcal{X}/B})-\frac{n\mu(\mathcal{X}_{b},\mathcal{L}_{b})}{n+1}c_{1}(\mathcal{L})^{n+1}\big{)}}{2V(\mathcal{X}_{b},\mathcal{L}_{b})}.$
2.1.1. Models over DVRs
We next discuss models, which will be the class of families of interest in the present paper. Let $R$ be a discrete valuation ring, endowed with a valuation $v$, with fraction field $K$ and residue field $\mathbb{C}$. Let $(X,L)$ be an $n$-dimensional normal $\mathbb{Q}$-polarized $K$-variety. This means that $(X,L)\to\operatorname{Spec}K$ is a flat family of $\mathbb{Q}$-polarized varieties of relative dimension $n$, where $L$ is a relatively ample $\mathbb{Q}$-line bundle.
Definition 2.2.
An $R$-model of $(X,L)$ is a flat normal projective relatively $\mathbb{Q}$-polarized integral scheme $(\mathcal{X},\mathcal{L})$ over $\operatorname{Spec}R$, together with an isomorphism $(\mathcal{X}_{K},\mathcal{L}_{K})\simeq(X,L)$ over $\operatorname{Spec}K$, where $(\mathcal{X}_{K},\mathcal{L}_{K})$ is the base change of $(\mathcal{X},\mathcal{L})$ over the generic point. Two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are said to be isomorphic if there exists an isomorphism over $\operatorname{Spec}R$ between them that extends the composition of isomorphisms $(\mathcal{X}_{K},\mathcal{L}_{K})\simeq(X,L)\simeq(\mathcal{X}_{K}^{\prime},\mathcal{L}_{K}^{\prime})$.
The CM line bundles of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are canonically isomorphic over the generic fiber, and using the identification with $(X,L)$, one may identify the difference of CM line bundles as a rational number, as follows: by functoriality of the Knudsen–Mumford expansion, the identification $(\mathcal{X}_{K},\mathcal{L}_{K})\simeq(X,L)\simeq(\mathcal{X}_{K}^{\prime},\mathcal{L}_{K}^{\prime})$ gives rise to a canonical isomorphism
$\big{(}\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\big{)}\big{|}_{\operatorname{Spec}K}\simeq\mathcal{O}_{\operatorname{Spec}K},$
of line bundles on $\operatorname{Spec}K$. This isomorphism may also be viewed as an isomorphism of $K$-vector spaces
$\varphi:H^{0}(\operatorname{Spec}K,\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))\to K,$
where the domain contains $\Lambda=H^{0}(\operatorname{Spec}R,\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))\cong R$.
Definition 2.3.
The difference of CM degrees of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ of $(X,L)$ is defined as the valuation
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=v(\gamma),$
where $\gamma$ is any non-zero element $K$ such that $\gamma\cdot R=\varphi(\Lambda)$. We write $\mathrm{CM}(\mathcal{X},\mathcal{L})\geq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, whenever $\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\geq 0$, and similarly for $\leq$ and strict inequalities.
Within K-stability, the idea of associating a number to two line bundles over the formal disk endowed with a fixed identification away from the special fiber goes back to Donaldson [Don12] (see also Wang [Wan12]). In the case $R=\mathbb{C}[\![t]\!]$, the number $v(\gamma)$ coincides with the unique largest integer $\nu$ such that $t^{-\nu}\varphi(s|_{\operatorname{Spec}\mathbb{C}(\!(t)\!)})$ extends to a regular function on $\operatorname{Spec}\mathbb{C}[\![t]\!]$.
One may describe the difference of CM degrees in terms of coefficients of Hilbert polynomials associated to the models: the composition of isomorphisms $\mathcal{X}^{\prime}_{K}\simeq X\simeq\mathcal{X}_{K}$ gives rise to a birational map $\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$. Choose a resolution of indeterminacies
(2.1) ${\mathcal{Y}}$${\mathcal{X}^{\prime}}$${\mathcal{X},}$$\scriptstyle{\mu^{\prime}}$$\scriptstyle{\mu}$
compatible with the morphisms to $\operatorname{Spec}R$. Write $\mu^{*}\mathcal{L}=\mu^{\prime*}\mathcal{L}^{\prime}+E-a\mathcal{Y}_{0}$ for an effective $\mathbb{Q}$-divisor $E$ supported in $\mathcal{Y}_{0}$ and a rational number $a>0$. This gives rise to an embedding
$H^{0}(\mathcal{Y},k\mathcal{M})\subset H^{0}(\mathcal{X},k\mathcal{L}),$
for $\mathcal{M}=\mu^{\prime*}\mathcal{L}^{\prime}-a\mathcal{Y}_{0}$ and sufficiently divisible $k$.
We may assume that $-E$ is $\mu$-relatively ample by replacing $(\mathcal{Y},\mathcal{M})$ with the ample model $(\operatorname{Proj}_{\mathcal{X}}(\bigoplus_{k\geq 0}\mu_{*}\mathcal{O}_{\mathcal{Y}}(k\mathcal{L})),\mathcal{O}(1))$ over $(\mathcal{X},\mathcal{L})$. There exists a closed subscheme $\mathcal{Z}$ in $\mathcal{X}$ supported in the special fiber such that the morphism $\mu:\mathcal{Y}\to\mathcal{X}$ is the blow-up of $\mathcal{X}$ along $\mathcal{Z}$ with exceptional divisor $rE$ for an integer $r\geq 1$. By [Laz17, Lemma 5.4.24], we obtain
$H^{i}(\mathcal{Y},k\mathcal{M})\simeq H^{i}(\mathcal{X},kr\mathcal{L}\otimes\mathcal{I}_{\mathcal{Z}}^{\otimes k}),$
for all $i\geq 0$ and sufficiently large $k$. In particular, the dimension of the quotient $H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})$ is finite and computed over the special fiber.
Theorem 2.4.
There exist rational numbers $b_{0}$ and $b_{1}$, such that
(2.2) $\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big{)}=b_{0}k^{n+1}+b_{1}k^{n}+O(k^{n-1})$
as $k\to+\infty$, for sufficiently divisible $k$. Moreover
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=\frac{b_{0}a_{1}-b_{1}a_{0}}{a_{0}^{2}},$
where $h^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})=h^{0}(\mathcal{X}_{0}^{\prime},k\mathcal{L}_{0}^{\prime})=a_{0}k^{n}+a_{1}k^{n-1}+O(k^{n-2})$.
In the case where the models are restrictions of flat families $(\overline{\mathcal{X}},\overline{\mathcal{L}})\to C$ and $(\overline{\mathcal{X^{\prime}}},\overline{\mathcal{L^{\prime}}})\to C$ for a smooth projective curve $C$, Hattori proved
$\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big{)}=\chi(\overline{\mathcal{X}},k\overline{\mathcal{L}})-\chi(\overline{\mathcal{X}^{\prime}},k(\overline{\mathcal{L}^{\prime}}+a\mathcal{X}_{0}^{\prime}))+O(k^{n-1}),$
[Hat24, Page 19]. It follows that Theorem 2.4 recovers Hattori’s definition of the difference of CM degrees of two models over a projective curve.
Proof.
We first claim that for all integers $i\geq 1$, the length of $H^{i}(\mathcal{Y},k\mathcal{M})$ as an $R$-module satisfies
(2.3) $\mathrm{length}_{R}H^{i}(\mathcal{Y},k\mathcal{M})=O(k^{n-1})$
as $k\to+\infty$. This was proven by Hattori [Hat24, Claim 1, Page 18] in the case where the family extends to a smooth projective curve. Indeed the same proof also applies for DVRs, as the Leray spectral sequence and Serre’s vanishing theorem hold over Noetherian rings.
We may assume that $R$ is regular by normalizing the base, pulling back the family and choosing an ample model; by functoriality of the Knudsen–Mumford expansion, the difference of CM degrees is unchanged, while Equation (2.2) is computed over the central fiber. In particular, one may choose a smooth proper curve $C$ and a point 00 in $C$, such that $R$ is isomorphic to the local ring $\mathcal{O}_{C,0}$ with $v=\operatorname{ord}_{0}$.
The identification of the models with $(X,L)$ over the generic fiber gives rise to an explicit trivialization of the vector bundle $E_{k}=\pi_{*}(k\mathcal{L})\otimes\pi^{\prime}_{*}(k\mathcal{L}^{\prime})^{\vee}$ away from zero, for sufficiently divisible $k$. Gluing this vector bundle on $\operatorname{Spec}K$ to the trivial bundle on the complement $C\setminus\{0\}$, we obtain vector bundle $\overline{E}_{k}$ on $C$ of rank $N_{k}=h^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})$, satisfying
$\chi(C,\overline{E}_{k})=\chi_{R}(\operatorname{Spec}R,E_{k})+\chi(C,\mathcal{O}^{N_{k}}_{C})-1,$
where $\chi_{R}(-)=\mathrm{rank}_{R}H^{0}(-)+\sum_{i>0}(-1)^{i}\mathrm{length}_{R}H^{i}(-)$, by the inclusion-exclusion formula for Euler characteristics.
On one hand, by construction, the degree of $\overline{E}_{k}$ over $C$ coincides with the degree of $E_{k}$, defined as
$\deg E_{k}:=v(\det\gamma),$
where the trivialization away from zero gives rise to a canonical isomorphism $\varphi:H^{0}(\operatorname{Spec}K,E_{k})\to K^{N_{k}}$ and $\gamma$ is any element in $\operatorname{GL}_{N_{k}}(K)$ such that $\gamma\cdot R^{N_{k}}=\varphi(\Lambda_{k})$, for the $R$-lattice $\Lambda_{k}=H^{0}(\operatorname{Spec}R,E_{k})$. On the other hand, by Riemann-Roch for projective curves, for large $k$, we have
$\deg\overline{E}_{k}$ $=\chi(C,\overline{E}_{k})-\chi(C,\mathcal{O}^{N_{k}}_{C}),$
$=\chi_{R}(\operatorname{Spec}R,\pi_{*}(k\mathcal{L}))-\chi_{R}(\operatorname{Spec}R,\pi^{\prime}_{*}(k\mathcal{L}^{\prime}))-1,$
$=\chi_{R}(\mathcal{X},k\mathcal{L})-\chi_{R}(\mathcal{Y},k\mathcal{M}^{\prime})+O(k^{n-1}),$
$=\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big{)}+O(k^{n-1}),$
using Equation 2.3 and $\chi_{R}(\operatorname{Spec}R,\pi_{*}(k\mathcal{L}))=\chi_{R}(\mathcal{X},k\mathcal{L})$ for large $k$.
Applying the Knudsen–Mumford expansion to $E_{k}$ and taking the degree, we obtain
(2.4) $b_{0}$ $=\frac{\deg(\lambda_{n+1}(\mathcal{X},\mathcal{L})-\lambda_{n+1}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))}{(n+1)!},$
$b_{1}$ $=-\frac{\deg(\lambda_{n}(\mathcal{X},\mathcal{L})-\lambda_{n}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))}{2n!},$
where the degree of the difference $\lambda_{i}(\mathcal{X},\mathcal{L})-\lambda_{i}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ is defined as for the difference of CM degrees. This yields the result. ∎
Remark 2.5.
With further hypotheses on the family $\mathcal{Y}\to\operatorname{Spec}R$, the difference of CM degrees may be defined in different ways. In the case where $\mathcal{Y}$ is normal and compactifies to a family over a smooth proper curve (still denoted $\mathcal{Y}$), we have
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=\frac{(\mu^{*}\mathcal{L}^{n}-\mathcal{M}^{n}).K_{\mathcal{Y}/R}-\frac{n\mu}{n+1}(\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1})}{2V}.$
If $\mathcal{Y}$ is smooth and the family $\mathcal{Y}\to\operatorname{Spec}R$ is the restriction of a family over a smooth (not necessarily proper) curve, the difference of CM degrees may also be viewed through Deligne pairings, as in this case the $\mathbb{Q}$-line bundles $\lambda_{i}$ in the Knudsen-Mumford expansion may be interpreted through Deligne pairings [PRS08]. Theorem 2.4 generalizes the formula for the Donaldson–Futaki invariant for test configurations and the difference of CM degrees for families over proper curves due to Odaka [Oda13] and Wang [Wan12].
Example 2.6.
Within K-stability, examples of models are ones induced by test configurations [Don02, Definition 2.1.1] (see also [DR24, Example 2.2]). A test configuration is a $\mathbb{C}^{*}$-equivariant degeneration $(\mathcal{Y},\mathcal{H})\to\mathbb{C}$ of a polarized variety $(Y,H)$ and its Donaldson–Futaki invariant is the difference of CM degrees
$\mathrm{DF}(\mathcal{Y},\mathcal{H})=\mathrm{CM}(\mathcal{Y}_{\mathbb{C}[\![t]\!]},\mathcal{H}_{\mathbb{C}[\![t]\!]})-\mathrm{CM}(Y\times\operatorname{Spec}\mathbb{C}[\![t]\!],H),$
between the restriction to the formal disk $\operatorname{Spec}\mathbb{C}[\![t]\!]$ and the trivial model.
2.1.2. Equivariant and twisted models
We introduce a subclass of models that are compatible with vertical torus actions on $(X,L)\to\operatorname{Spec}K$. Let $T$ be a $K$-torus in $\mathrm{Aut}_{K}(X,L)$, meaning $T\cong\mathbb{G}_{m,K}^{d}$ for some integer $d\geq 0$.
Definition 2.7.
A model $(\mathcal{X},\mathcal{L})$ of $(X,L)$ is called $T$-equivariant if it admits an $R$-torus $T_{R}$ in $\mathrm{Aut}_{R}(\mathcal{X},\mathcal{L})$ such that $T_{R}|_{(\mathcal{X}_{K},\mathcal{L}_{K})}=\varphi^{*}T$, where $\varphi:(\mathcal{X}_{K},\mathcal{L}_{K})\to(X,L)$ is the designated identification associated to the model $(\mathcal{X},\mathcal{L})$.
This is a generalization of invariant test configurations due to Székelyhidi [Szé07, Definition 2.1]. As the generic fiber $\mathcal{X}_{K}$ is dense in $\mathcal{X}$, if such a torus $T_{R}$ exists, it must be unique.
Choose a uniformizer $t$ in the discrete valuation ring, that is, an element in $R$ such that its valuation $v(t)$ generates the valuation group $v(K^{*})$ in $\mathbb{R}$; (in the case $R=\mathbb{C}[\![t]\!]$, the element $t$ itself is a uniformizer). Let $(\mathcal{X},\mathcal{L})$ be a model of $(X,L)$ and let $\beta$ be a vertical $\mathbb{G}_{m,K}$-action on $(X,L)$. The choice of uniformizer $t$ gives rise to an automorphism $\beta(t):(X,L)\to(X,L)$.
Definition 2.8.
The twisted model of $(\mathcal{X},\mathcal{L})$ by $\beta$ is defined to be the model $(\mathcal{X}_{\beta},\mathcal{L}_{\beta}):=(\mathcal{X},\mathcal{L})$ whose identification with $(X,L)$ has been replaced by $\varphi_{\beta}:=\beta(t)\circ\varphi$.
The definition of twisted models extends the notion of twisted test configurations introduced by Székelyhidi [Szé07], and further developed by Li [Li22]. Note that, unlike test configurations, models do not need to be equivariant in order for them to be twisted.
[…]
One may also twist by rational one-parameter subgroups $\xi$ of $T$; that is, elements of the vector space $\mathrm{Hom}(\mathbb{G}_{m,K},\operatorname{Aut}_{K}(X,L))\otimes\mathbb{Q}$. If $\xi=r^{-1}\beta$ for an integer $r\geq 1$ and a one-parameter subgroup $\beta$ of $T$, then we define the twist of $(\mathcal{X},\mathcal{L})$ by $\xi$ as the twist
$(\mathcal{X}_{\xi},\mathcal{L}_{\xi}):=(\mathcal{X},r^{-1}\mathcal{L}_{r\beta}).$
This means that the identification over $\operatorname{Spec}K$ takes the form
$(\mathcal{X}_{\xi,K},\mathcal{L}_{\xi,K})=(\mathcal{X},r^{-1}\mathcal{L})\simeq(X,r^{-1}L)\xrightarrow{r\beta(t)}(X,L).$
2.2. Invariant, twisted and good filtrations.
In this section we recall the notions of filtrations and Okounkov bodies of complex polarized varieties, following the treatment in [Szé15] (up to a change in sign) and give the definitions of invariant, twisted and good filtrations, introduced by Li [Li22] and Hattori [Hat24].
Let $(Y,H)$ be an $n$-dimensional complex polarized variety with ring of sections $S=\bigoplus_{k\geq 0}S_{k}$, where $S_{k}=H^{0}(Y,kH)$.
Definition 2.9.
A (unitary, decreasing, multiplicative, linearly bounded) $\mathbb{Z}$-filtration of $(Y,H)$ is a collection $\mathcal{F}=(\mathcal{F}^{\lambda}S_{k})_{\lambda,k}$ of linear subspaces $\mathcal{F}^{\lambda}S_{k}$ in $S_{k}$ for $\lambda$ in $\mathbb{Z}$ and $k$ in $\mathbb{Z}_{\geq 0}$ such that for all integers $\lambda\leq\mu$ and and all integers $k,l\geq 0$, we have
  1. (i)
    $1\in\mathcal{F}^{\lambda}S_{0}$,
  2. (ii)
    $\mathcal{F}^{\lambda}S_{k}\supset\mathcal{F}^{\mu}S_{k}$,
  3. (iii)
    $\mathcal{F}^{\lambda}S_{k}\cdot\mathcal{F}^{\mu}S_{l}\subset\mathcal{F}^{\lambda+\mu}S_{k+l}$,
  4. (iv)
    $\mathcal{F}^{\lambda}S_{k}=S_{k}$ for all $\lambda<-Ck$ and $\mathcal{F}^{\lambda}S_{k}=0$ for all $\lambda\geq-Ck$,
for some constant $C>0$, independent from $\lambda$ and $k$. The $p$th-order weight function of $\mathcal{F}$ is defined to be the function $w_{\mathcal{F}}^{p}:\mathbb{Z}\to\mathbb{Z}$, given by
$w_{\mathcal{F}}^{p}(k)=\sum_{\lambda\in\mathbb{Z}}\lambda^{p}\dim(\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k}).$
We write $w_{\mathcal{F}}=w_{\mathcal{F}}^{1}$. The $L^{2}$-norm of $\mathcal{F}$ is given by
$||\mathcal{F}||^{2}=\frac{c_{0}a_{0}-b_{0}^{2}}{a_{0}^{2}},$
where $a_{0}$, $b_{0}$ and $c_{0}$ are real numbers coming from the expansions
(2.5) $\dim S_{k}$ $=a_{0}k^{n}+O(k^{n-1}),$
$w_{\mathcal{F}}(k)$ $=b_{0}k^{n+1}+o(k^{n+1}),$
$w_{\mathcal{F}}^{2}(k)$ $=c_{0}k^{n+2}+o(k^{n+2}),$
as $k\to+\infty$. We call $\mathcal{F}$ finitely generated if the $\mathbb{C}[t]$-algebra $\bigoplus_{k,\lambda}\mathcal{F}^{\lambda}S_{k}\cdot t^{-\lambda}$ is finitely generated.
The numbers $b_{0}$ and $c_{0}$ exist as the filtration is linearly bounded.
Example 2.10 ([Nys12, Definition 6.2]).
An important example of filtrations are the ones induced by test configurations. In the case of a product test configuration $(Y\times\mathbb{C},H)_{\beta}$ induced by a $\mathbb{C}^{*}$-action $\beta$ on $(Y,H)$, the associated filtration $\mathcal{F}_{(Y\times\mathbb{C},H)_{\beta}}$ takes the form
($\mathcal{F}^{\lambda}_{(Y\times\mathbb{C},H)_{\beta}}S_{k}=\bigoplus_{\mu\geq\lambda}(S_{k})_{\mu},$
where $S_{k}=\bigoplus_{\mu\in\mathbb{Z}}(S_{k})_{\mu}$ denotes the weight-space decomposition with respect to the $\mathbb{C}^{*}$-action induced by $\beta$. In particular, anything that is defined for filtrations is also defined for $\mathbb{C}^{*}$-actions through the associated filtration.
Example 2.11 ([Szé15, Section 3.2]).
Any filtration $\mathcal{F}$ may be approximated by finitely generated filtrations $\mathcal{F}_{(r)}$ of $(Y,rH)$ for $r\geq 1$, given by
($\mathcal{F}^{\lambda}_{(r)}S_{kr}=\{\sigma\;|\;\sigma\cdot t^{-\lambda}\in\mathcal{S}_{kr}(\mathcal{F})\},$
where $\mathcal{S}_{k}(\mathcal{F})\subset S[t]$ is the $\mathbb{C}[t]$-subalgebra generated by $\bigoplus_{\lambda}\mathcal{F}^{\lambda}S_{k}\cdot t^{-\lambda}$. We call the sequence of filtrations $(\mathcal{F}_{(r)})$ the tautological approximation of $\mathcal{F}$.
2.2.1. Okounkov bodies and filtrations.
Okounkov bodies [Oko96] play an important role in capturing asymptotic data of ample line bundles and filtrations. For this section, assume that $(Y,H)$ is smooth.
Given local coordinates $(U,z)$ centered around a point $p$ in $Y$ and a global section $\sigma$ of $L$ which does not vanish at $p$, we define a convex subset
(2.6) $P=\overline{\bigcup_{k\geq 1}\frac{P_{k}}{k}}\subset\mathbb{R}^{n},$
where $P_{k}=\{\nu_{\sigma}(f)\;|\;f\in S_{k}\}\subset\mathbb{Z}^{n}$ and $\nu_{\sigma}(f)$ is the multi-degree of the lowest-order term of the power series $q$ in $\mathbb{Z}[\![z_{1},\dots,z_{n}]\!]$ such that $f=\sigma^{k}q$ on $U$ with respect to the lexicographic order. We call $P$ the Okounkov body of $(Y,H)$ centered around $p$, induced by the data $(U,z,\sigma)$. Important properties are $|P_{k}|=\dim S_{k}$ and $\mathrm{vol}(P)=n!\mathrm{vol}(L)$. From [Szé15, Lemma 6], if follows that for sufficiently large integers $k$ and $r$, we have $\Delta_{k-1}\cap\mathbb{Z}^{n}\subset P_{kr}$, where $\Delta_{l}=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}\;|\;x_{i}\geq 0\text{ and }\sum_{i=1}^{n}x_{i}\leq l\}$ denotes the closed solid $n$-simplex of side length $l>0$ In particular, we have $\Delta_{1/r}\subset P$.
Let $\mathcal{F}$ be a filtration of $(Y,H)$.
Definition 2.12 ([Nys12, Section 3], [Szé15, Equation 13]).
The $k$th-lattice function and the concave transform of $\mathcal{F}$ with respect to $P$ are the functions $g_{\mathcal{F},k}:P_{k}\to\mathbb{Z}$ and $G_{\mathcal{F}}:P\to\mathbb{R}\cup\{-\infty\}$, given by
$g_{\mathcal{F},k}(a)$ $=\max\{\lambda\in\mathbb{Z}\;|\;a\in P_{k}(\tfrac{\lambda}{k},\mathcal{F})\},$
$G_{\mathcal{F}}(x)$ $=\sup\{s\in\mathbb{R}\;|\;x\in P(s,\mathcal{F})\},$
where $P_{k}(s,\mathcal{F})=\nu_{\sigma}(\mathcal{F}^{\lceil sk\rceil}S_{k})\subset P_{k}$ and $P(s,\mathcal{F})=\overline{\bigcup_{k\geq 1}\frac{P_{k}(s,\mathcal{F})}{k}}\subset P$.
The function $G_{\mathcal{F}}$ is concave and upper semi-continuous on $P$ and is finite and continuous on the interior of $P$. A key property is that one may express the leading coefficient of the $p$th-order weight function of $\mathcal{F}$ as
(2.7) $w_{\mathcal{F}}^{p}(k)$ $=\Big{(}\int_{P}G_{\mathcal{F}}^{p}\Big{)}k^{n+p}+o(k^{n+p}),$
as $k\to+\infty$, integrating with respect to the Lebesgue measure (as we do throughout our work). In particular, the $L^{2}$-norm of $\mathcal{F}$ may be written as
$||\mathcal{F}||^{2}=\frac{1}{\mathrm{vol}(P)}\int_{P}(G_{\mathcal{F}}-\overline{G}_{\mathcal{F}})^{2},$
where $\overline{G}_{\mathcal{F}}=\frac{1}{\mathrm{vol}(P)}\int_{P}G_{\mathcal{F}}$ denotes the average of $G_{\mathcal{F}}$ over $P$.
2.2.2. Invariant and twisted filtrations.
We introduce a subclass of filtrations which serves as a generalization of invariant test configurations [Szé07, Definition 2.1]. Let $\beta$ be a $\mathbb{C}^{*}$-action on $(Y,H)$.
Definition 2.13 ([Li22, Section 3.2]).
A filtration $\mathcal{F}$ of $(Y,H)$ is called $\beta$-invariant if each of its filtered pieces $\mathcal{F}^{\lambda}S_{k}$ are invariant under the $\mathbb{C}^{*}$-action on $S_{k}$ induced by $\beta$.
A $\beta$-invariant filtration $\mathcal{F}$ gives rise to weight-space decompositions
$\mathcal{F}^{\lambda}S_{k}$ $=\bigoplus_{\mu\in\mathbb{Z}}\big{(}\mathcal{F}^{\lambda}S_{k}\big{)}_{\mu},$
$\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k}$ $=\bigoplus_{\mu\in\mathbb{Z}}\big{(}\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k}\big{)}_{\mu},$
admitting natural identifications $(\mathcal{F}^{\lambda}S_{k})_{\mu}/(\mathcal{F}^{\lambda+1}S_{k})_{\mu}\simeq(\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}$.
Definition 2.14.
The mixed square-weight function of a $\beta$-invariant filtration $\mathcal{F}$ and $\beta$ is the function $w_{\mathcal{F},\beta}^{2}:\mathbb{Z}\to\mathbb{Z}$, given by
$w^{2}_{\mathcal{F},\beta}(k)=\sum_{\lambda,\mu\in\mathbb{Z}}\lambda\mu\dim\big{(}(\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}\big{)},$
The $L^{2}$-inner product of $\mathcal{F}$ and $\beta$ is defined as
$\langle\mathcal{F},\beta\rangle=\frac{c_{0}^{\prime}a_{0}-b_{0}b_{0}^{\prime}}{a_{0}^{2}},$
where $a_{0}$ and $b_{0}$ are defined as in Equation (2.5), and, as $k\to+\infty$,
$w_{\beta}(k)$ $=b_{0}^{\prime}k^{n+1}+O(k^{n}),$
$w^{2}_{\mathcal{F},\beta}(k)$ $=c_{0}^{\prime}k^{n+2}+o(k^{n+2}).$
The mixed square-weight and the inner product naturally extend the usual definitions for commuting $\mathbb{C}^{*}$-actions on $(Y,H)$ in [Szé07].
Definition 2.15 ([Li22, Definition 3.6]).
The twisted filtration of a $\beta$-invariant filtration $\mathcal{F}$ by $\beta$ is the filtration $\mathcal{F}_{\beta}$ of $(Y,H)$, defined by
$\mathcal{F}_{\beta}^{\lambda}S_{k}=\bigoplus_{\lambda_{1}+\lambda_{2}=\lambda}\big{(}\mathcal{F}^{\lambda_{1}}S_{k}\big{)}_{\lambda_{2}}.$
The weight pieces of the twisted filtration are $(\mathcal{F}_{\beta}^{\lambda}S_{k})_{\mu}=(\mathcal{F}^{\lambda-\mu}S_{k})_{\mu}$.
Lemma 2.16.
The first- and second-order weight functions of $\mathcal{F}_{\beta}$ satisfy
$w_{\mathcal{F}_{\beta}}$ $=w_{\mathcal{F}}+w_{\beta},$
$w_{\mathcal{F}_{\beta}}^{2}$ $=w_{\mathcal{F}}^{2}+2w^{2}_{\mathcal{F},\beta}+w^{2}_{\beta}.$
In particular, we have
$||\mathcal{F}_{\beta}||^{2}=||\mathcal{F}||^{2}+2\langle\mathcal{F},\beta\rangle+||\beta||^{2}.$
In the case of test configurations, this is proven using that the infinitesimal generator of a twisted test configuration is the sum of infinitesimal generators of the test configuration and the vertical $\mathbb{C}^{*}$-action. In the proof for filtrations, this is hidden behind the definition of a twisted filtration which is chosen in such a way that it matches the definition of a twisted test configuration.
Proof.
Writing $S_{k}=\bigoplus_{\mu\in\mathbb{Z}}(S_{k})_{\mu}$ for the weight-space decomposition with respect to the $\beta$-action, we compute directly:
$w_{\mathcal{F}_{\beta}}(k)$ $=\sum_{\lambda_{1},\lambda_{2}\in\mathbb{Z}}(\lambda_{1}+\lambda_{2})\dim\big{(}(\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{1}+1}S_{k})_{\lambda_{2}}\big{)},$
$=\sum_{\lambda_{1}\in\mathbb{Z}}\lambda_{1}\dim\Big{(}\bigoplus_{\lambda_{2}\in\mathbb{Z}}\big{(}\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{1}+1}S_{k}\big{)}_{\lambda_{2}}\Big{)}$
$\qquad+\sum_{\lambda_{2}\in\mathbb{Z}}\lambda_{2}\dim\Big{(}\bigoplus_{\lambda_{1}\in\mathbb{Z}}\big{(}\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{1}+1}S_{k}\big{)}_{\lambda_{2}}\Big{)},$
$=\sum_{\lambda_{1}\in\mathbb{Z}}\lambda_{1}\dim\big{(}\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{2}+1}S_{k}\big{)}+\sum_{\lambda_{2}\in\mathbb{Z}}\lambda_{2}\dim\big{(}(S_{k})_{\lambda_{2}}\big{)},$
$=w_{\mathcal{F}}(k)+w_{\beta}(k).$
using that there is an isomorphism $\bigoplus_{\lambda_{1}\in\mathbb{Z}}(\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{1}+1}S_{k})_{\lambda_{2}}\cong(S_{k})_{\lambda_{2}}$. Following a similar computation for the square-weight, we obtain
$w^{2}_{\mathcal{F}_{\beta}}(k)$ $=\sum_{\lambda_{1}\in\mathbb{Z}}\lambda_{1}^{2}\dim(\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{2}+1}S_{k})$
$\quad+2\sum_{\lambda_{1},\lambda_{2}\in\mathbb{Z}}\lambda_{1}\lambda_{2}\dim\big{(}(\mathcal{F}^{\lambda_{1}}S_{k}/\mathcal{F}^{\lambda_{1}+1}S_{k})_{\lambda_{2}}\big{)}$
$\quad+\sum_{\lambda_{2}\in\mathbb{Z}}\lambda_{2}^{2}\dim\big{(}(S_{k})_{\lambda_{2}}\big{)},$
$=w_{\mathcal{F}}^{2}(k)+2w^{2}_{\mathcal{F},\beta}(k)+w^{2}_{\beta}(k),$
as desired. ∎
A modification of [Nys12, Lemma 3.3] and [BC11, Theorem A] gives rise to the following relation and asymptotics between the weight pieces, lattice functions and concave transforms:
Lemma 2.17.
For any function $T:\mathbb{Z}^{2}\to\mathbb{R}$, we have
$\sum_{\lambda,\mu\in\mathbb{Z}}T(\lambda,\mu)\dim((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu})=\sum_{a\in P_{k}}T(g_{\mathcal{F},k}(a),g_{\beta,k}(a)).$
Moreover, for any continuous function $T:\mathbb{R}^{2}\to\mathbb{R}$ such that $T\circ(G_{\mathcal{F}},G_{\beta})$ is integrable on $P$ with respect to the Lebesgue measure, we have
$\lim_{k\to\infty}\frac{\sum_{a\in P_{k}}T\big{(}\frac{g_{\mathcal{F},k}(a)}{k},\frac{g_{\beta,k}(a)}{k}\big{)}}{k^{n}}=\int_{P}T\circ(G_{\mathcal{F}},G_{\beta}).$
Proof.
From [Nys12, Equation 1] and the definition of the lattice functions, we have $\dim((\mathcal{F}^{\lambda}R_{k})_{\mu})=\sum_{a\in P_{k}}\delta_{a}(g_{\mathcal{F},k}\geq\lambda)\delta_{a}(g_{\beta,k}=\mu)$ where $\delta_{a}$ is the Dirac measure on $P_{k}$ concentrated at a point $a$. Thus,
$\sum_{\lambda,\mu\in\mathbb{Z}}$ $T(\lambda,\mu)\dim((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}),$
$=\sum_{\lambda,\mu\in\mathbb{Z}}T(\lambda,\mu)\sum_{a\in P_{k}}\big{(}\delta_{a}(g_{\mathcal{F},k}\geq\lambda)-\delta_{a}(g_{\mathcal{F},k}\geq\lambda+1)\big{)}\delta_{a}(g_{\beta,k}=\mu),$
$=\sum_{a\in P_{k}}\sum_{\lambda,\mu\in\mathbb{Z}}T(\lambda,\mu)\delta_{a}(g_{\mathcal{F},k}=\lambda)\delta_{a}(g_{\beta,k}=\mu),$
$=\sum_{a\in P_{k}}T(g_{\mathcal{F},k}(a),g_{\beta,k}(a)).$
The subset $P(t,\mathcal{F})$ in $P$ is the Okounkov subbody of the graded subalgebra $S(t,\mathcal{F})=\bigoplus_{k\geq 0}S_{k}(t,\mathcal{F})$ for $S_{k}(t,\mathcal{F})=\mathcal{F}^{\lceil tk\rceil}S_{k}$. Hence, by [BC11, Equation 1.8] and [LM09, Theorem 2.13], in the case where $T$ is smooth, we have
$\int_{P}T\circ(G_{\mathcal{F}},G_{\beta})$ $=\int_{\mathbb{R}^{2}}DT(t,s)(\mathrm{vol}(G_{\mathcal{F}}\geq t),\mathrm{vol}(G_{\beta}\geq s)),$
$=\int_{\mathbb{R}^{2}}DT(t,s)(\mathrm{vol}(P(t,\mathcal{F})),\mathrm{vol}(P(t,\beta))),$
$=\lim_{k\to\infty}\int_{\mathbb{R}^{2}}DT(t,s)\Big{(}\frac{\dim S_{k}(t,\mathcal{F})}{k^{n}},\frac{\dim S_{k}(t,\beta)}{k^{n}}\Big{)},$
$=\lim_{k\to\infty}\sum_{\lambda,\mu\in\mathbb{Z}}T\Big{(}\frac{\lambda}{k},\frac{\mu}{k}\Big{)}\frac{\dim((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu})}{k^{n}},$
$=\lim_{k\to\infty}\frac{\sum_{a\in P_{k}}T\big{(}\frac{g_{\mathcal{F},k}(a)}{k},\frac{g_{\beta,k}(a)}{k}\big{)}}{k^{n}}.$
where we may change the limit with the integral as the volume of $P(t,\mathcal{F})$ is bounded uniformly by the volume of $P$. The result for general continuous functions follows by density. ∎
In particular, the mixed square-weight function between $\mathcal{F}$ and $\beta$ satisfies
(2.8) $w_{\mathcal{F},\beta}^{2}(k)$ $=\Big{(}\int_{P}G_{\mathcal{F}}G_{\beta}\Big{)}k^{n+2}+o(k^{n+2}),$
as $k\to+\infty$, and the $L^{2}$-inner product of $\mathcal{F}$ and $\beta$ may be expressed as
$\langle\mathcal{F},\beta\rangle=\frac{1}{\mathrm{vol}(P)}\int_{P}(G_{\mathcal{F}}-\overline{G}_{\mathcal{F}})(G_{\beta}-\overline{G}_{\beta}).$
2.2.3. Torus-invariant filtrations
We conclude the section with the notion of torus-invariant filtrations and the associated reduced $L^{2}$-norm. Let $T$ be a complex torus of rank $d$ in the automorphism group $\mathrm{Aut}(Y,H)$, and write
$N_{\mathbb{Z}}(T)=\mathrm{Hom}(\mathbb{C}^{*},T).$
The group $N_{\mathbb{Z}}(T)$ is free abelian of rank $d$ whose elements are $\mathbb{C}^{*}$-actions on $(Y,H)$. We write the group operation in $N_{\mathbb{Z}}(T)$ additively.
Definition 2.18.
A filtration $\mathcal{F}$ of $(Y,H)$ is called $T$-invariant, if all of its filtration pieces $\mathcal{F}^{\lambda}S_{k}$ are invariant under the natural linear $T$-action on $S_{k}$.
A filtration $\mathcal{F}$ is $T$-invariant if and only if it is $\beta$-invariant for all $\beta$ in $N_{\mathbb{Z}}(T)$. The mixed-square weight $\langle\mathcal{F},\beta\rangle$ extends bilinearly to the real vector space $N_{\mathbb{R}}(T)=N_{\mathbb{Z}}(T)\otimes\mathbb{R}$. This allows us to make the following definition:
Definition 2.19.
The reduced norm of a $T$-invariant filtration $\mathcal{F}$ of $(Y,H)$ is given by the infimum
$||\mathcal{F}||_{T}^{2}=\inf_{\xi\in N_{\mathbb{R}}(T)}(||\mathcal{F}||^{2}+2\langle\mathcal{F},\xi\rangle+||\xi||^{2}),$
where $||\xi||^{2}=\langle\xi,\xi\rangle$ is the natural extension of the norm to $N_{\mathbb{R}}(T)$.
In light of Lemma 2.16, the expression in the infimum may be thought of as the norm of the twist $\mathcal{F}_{\xi}$ by a vector $\xi$ as an $\mathbb{R}$-filtration (see [Li22, Definition 2.28]). The infimum is actually a minimum:
Lemma 2.20.
We have
$||\mathcal{F}||_{T}^{2}=||\mathcal{F}||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle+||\xi(\mathcal{F})||^{2},$
for $\xi(\mathcal{F})=\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\beta_{i}$ where $(\beta_{i})$ is any $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$.
Proof.
For any vector $\zeta$ in $N_{\mathbb{R}}(T)$, we have $\langle\xi(\mathcal{F}),\zeta\rangle=\langle\mathcal{F},\zeta\rangle$, implying
$||\xi(\mathcal{F})+\zeta||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})+\zeta\rangle=||\xi(\mathcal{F})||^{2}+||\zeta||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle.$
This is bounded from below by $||\xi(\mathcal{F})||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle$, as desired. ∎
2.2.4. Good filtrations
An issue with non-finitely-generated filtrations is that, in general, we have no control over subleading coefficients of the weight function. This makes it difficult to define the polynomial Donaldson–Futaki invariant of general filtrations, so we introduce a subclass of filtrations for which we have control over the weight function as a polynomial.
Definition 2.21 ([Hat24, Definition 2.14]).
A filtration $\mathcal{F}$ of $(Y,H)$ is called good if there is a polynomial $b(t)$ over $\mathbb{Q}$ such that $w_{\mathcal{F}}(k)=b(k)$ for sufficiently large integers $k$. The Donaldson–Futaki invariant of a good filtration $\mathcal{F}$ is defined as
$\mathrm{DF}(\mathcal{F})=\frac{b_{0}a_{1}-b_{1}a_{0}}{a_{0}^{2}},$
where we have written $h^{0}(Y,kH)=a_{0}k^{n}+a_{1}k^{n-1}+O(k^{n-2})$ as $k\to+\infty$.
Finitely generated filtrations are good and the definition the Donaldson–Futaki invariant coincides with the usual definitions for test configurations. Using the weight formula in Lemma 2.16, we obtain the following:
Lemma 2.22.
If $\mathcal{F}$ is good and invariant with respect to a $\mathbb{C}^{*}$-action $\beta$ on $(Y,H)$, then
$\mathrm{DF}(\mathcal{F}_{\beta})=\mathrm{DF}(\mathcal{F})+\mathrm{Fut}(\beta),$
where $\mathrm{Fut}(\beta)=\mathrm{DF}(\beta)$ is the Futaki invariant of $\beta$ [Fut83].
Remark 2.23 ([Hat24, Remark 2.21]).
The definition of the Donaldson-Futaki invariant of a good filtration in principle differs from the definition in [Szé15, Definition 3] in the sense that
$\mathrm{DF}(\mathcal{F})\leq\liminf_{r\to+\infty}\mathrm{DF}(\mathcal{F}_{(r)}),$
for good filtrations $\mathcal{F}$. It is not known whether this is an equality in general; this is a problem about the convergence of subleading coefficients of the weight function, and is related to the non-Archimedean entropy approximation conjecture [BJ18, Conjecture 2.5] and [Li22, Conjecture 1.8].
3. Relative stability of extremal polarized manifolds
Let $(Y,H)$ be a smooth polarized variety and let $T$ in $\mathrm{Aut}(Y,H)$ be a maximal complex torus of rank $d$. The goal of this section is to prove that if $(Y,H)$ is extremal, then it is relatively K-polystable with respect to good filtrations. This is second step outlined in Section 1.3 used in the proof of 1.2. We start by introducing the relevant numerical invariant.
Definition 3.1.
Let $\mathcal{F}$ be a $T$-invariant good filtration of $(Y,H)$. The relative Donaldson–Futaki invariant of $\mathcal{F}$ of $(Y,H)$ is defined as
$\mathrm{DF}_{T}(\mathcal{F})=\mathrm{DF}(\mathcal{F})-\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\mathrm{Fut}(\beta_{i}),$
where $(\beta_{i})$ is any choice of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)=\mathrm{Hom}(\mathbb{C}^{*},T)$.
This extends the definition of the relative Donaldson–Futaki invariant for test configurations [Szé07, Equation 1.1].
We main goal of this section is the following:
Theorem 3.2.
If $(Y,H)$ is extremal, then for any $T$-invariant good filtration $\mathcal{F}$ of $(Y,H)$, we have
$\mathrm{DF}_{T}(\mathcal{F})\geq 0,$
where the inequality is strict if and only if $||\mathcal{F}||_{T}>0$.
This is a simultaneous generalization of [Hat24, Theorem 2.18], [Szé15, Theorem 10] and [SS11, Theorem 4] allowing automorphisms and filtrations. In the case where $(Y,H)$ admits a constant scalar curvature Kähler (cscK) metric, the Futaki invariant in Lemma 2.22 vanishes, giving rise to the following:
Corollary 3.3.
If $(Y,H)$ admits a cscK metric, then for any $T$-invariant good filtration $\mathcal{F}$ of $(Y,H)$, we have
$\mathrm{DF}(\mathcal{F})\geq 0,$
where the inequality is strict if and only if $||\mathcal{F}||_{T}>0$.
The general strategy for proving Theorem 3.2 involves first showing that the relative Donaldson–Futaki invariant is bounded from below by the relative asymptotic Chow weight (Proposition 3.15) and proving that extremal manifolds are relatively asymptotically filtration Chow stable (Theorem 3.16). For the latter, we follow a method originating with Stoppa [Sto07] by first establishing a blowup formula for the relative asymptotic Chow weight, then finding a destabilizing point to blowup and then applying results by Arezzo–Pacard–Singer [APS11, Theorem 2.1] and Seyyedali [Sey17, Corollary 1.2] on relative asymptotic Chow semistability of blowups of extremal manifolds (see also [Mab18, ST21, Has21]) to obtain a contradiction.
3.1. Székelyhidi’s formula
We start with recalling Szkékelyhidi’s blowup formula for the Chow weight of general filtrations in [Szé15, Theorem 10]. Given a filtration $\mathcal{F}$ of $(Y,H)$, the $r$th-Chow weight of $\mathcal{F}$ is defined as
$\mathrm{Chow}_{r}(\mathcal{F})=\frac{rb_{0}}{a_{0}}-\frac{w_{\mathcal{F}}(r)}{h^{0}(Y,rH)},$
where $a_{0}$ and $b_{0}$ are as in Equation (2.5). Choose a point $p$ in $Y$ and write $(\widehat{Y},\widehat{H}_{m})=(\mathrm{Bl}_{p}Y,mH-E)$ for the blowup with exceptional divisor $E$, which is a polarized variety for sufficiently large integers $m$. Given a filtration $\mathcal{F}$ of $(Y,H)$, we define a new filtration $\widehat{\mathcal{F}}$ of $(\widehat{Y},\widehat{H}_{m})$ by
(3.1) ^$\widehat{\mathcal{F}}^{\lambda}H^{0}\big{(}\widehat{Y},k\widehat{H}_{m}\big{)}=\{\sigma\;|\;i(\sigma)\in\mathcal{F}^{\lambda}H^{0}(Y,kmH)\},$
where $i:H^{0}(\widehat{Y},k\widehat{H}_{m})\hookrightarrow H^{0}(Y,kmH)$ is the natural embedding of sections. This filtration generalizes the construction of a test configuration on blowups of Stoppa [Sto07, Section 2].
From now on, we will assume $n\geq 2$. We will refer to the case $n\leq 1$ in Remark 3.17 when discussing the stability results. Székelyhidi’s formula is the following asymptotic bound on the Chow weight on the blowup:
Theorem 3.4 ([Szé15, Page 18]).
Let $P$ be an Okounkov body of $(Y,H)$ centered at $p$. Then
$\mathrm{Chow}_{r}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{mr}(\mathcal{F})+\Big{(}\nu_{m}(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big{)}\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0}),$
as $r\to+\infty$ and subsequently $m\to+\infty$, where
(3.2) $\nu_{m}(p,\mathcal{F})=\frac{(n-2)!a_{0}}{m^{1-n}}\bigg{(}\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}}-\frac{1}{\sqrt{n}}\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}}\bigg{)}.$
Here, $\partial_{0}\Delta_{l}$ denotes the union of boundary faces meeting in the origin the solid closed $n$-simplex $\Delta_{l}\subset\mathbb{R}^{n}$ of length $l$ and $\partial_{1}\Delta_{l}$ denotes the remaining face. The value in Equation (3.2) depends on the choice of Okounkov body; however, we always have $\nu_{m}(\mathcal{F},p)=O(m^{0})$ as $m\to+\infty$.
3.2. A relative formula.
We next establish the a relative blowup formula, extending Theorem 3.4, needed in the proof of Theorem 3.16. Given a $\mathbb{C}^{*}$-action $\beta$ on $(Y,H)$ and a $\beta$-invariant filtration $\mathcal{F}$ be $(Y,H)$, we define
(3.3) $\langle\mathcal{F},\beta\rangle_{r}=\frac{w^{2}_{\mathcal{F},\beta}(r)h^{0}(Y,rH)-w_{\mathcal{F}}(r)w_{\beta}(r)}{r^{2}h^{0}(Y,rH)^{2}}.$
By definition of the $L^{2}$-inner product, we have
$\langle\mathcal{F},\beta\rangle=\lim_{r\to+\infty}\langle\mathcal{F},\beta\rangle_{r}.$
Assume now that $p$ is fixed by $\beta$. In this case, the $\mathbb{C}^{*}$-action $\beta$ naturally lifts to a $\mathbb{C}^{*}$-action $\widehat{\beta}$ on the blowup $(\widehat{Y},\widehat{H}_{m})$ leaving $\widehat{\mathcal{F}}$ invariant.
Proposition 3.5.
As $r\to+\infty$ and $m\to+\infty$, we have
^$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle_{r}=\langle\mathcal{F},\beta\rangle_{mr}m^{2}+o(m^{2-n}r^{0}).$
Proof.
By asymptotic Riemann-Roch we have
$h^{0}(\widehat{Y},r\widehat{H}_{m})=h^{0}(Y,mrH)+O(m^{0}r^{n})$
as $r\to+\infty$ and $m\to+\infty$.
Given a choice of Okounkov body $P$ for $(Y,H)$ centered at $p$ constructed out of data $(U,z,\sigma)$, one obtains an Okounkov body
$\widehat{P}=\overline{mP\setminus\Delta_{1}}$
for $(\widehat{Y},\widehat{H}_{m})$ constructed out of lifted data $(\widehat{U},\widehat{z},\widehat{\sigma^{m}})$. A direct computation shows that $G_{\widehat{\mathcal{F}}}(x)=mG_{\mathcal{F}}(\frac{x}{m})$ for all $x$ in $\widehat{P}$. Using this and Equation (2.7), we have
$w_{\widehat{\mathcal{F}}}(r)$ $=\bigg{(}\int_{mP}G_{\mathcal{F}}\bigg{)}m^{n+1}r^{n+1}-\bigg{(}\int_{\Delta_{\frac{1}{m}}}G_{\mathcal{F}}\bigg{)}m^{n+1}r^{n+1}+o(r^{n+1}),$
$=w_{\mathcal{F}}(mr)+O(mr^{n+1}).$
Similarly, using Equation (2.8), we have
$w^{2}_{\widehat{\mathcal{F}},\widehat{\beta}}(r)$ $=\bigg{(}\int_{P}G_{\mathcal{F}}G_{\beta}\bigg{)}m^{n+2}r^{n+2}-\bigg{(}\int_{\Delta_{\frac{1}{m}}}G_{\mathcal{F}}G_{\beta}\bigg{)}m^{n+2}r^{n+2}+o(r^{n+2}),$
$=w^{2}_{\mathcal{F},\beta}(mr)+O(m^{2}r^{n+2}).$
Putting these expansions in Equation (3.3) and taking the limits in the right order yields the result. ∎
Remark 3.6.
Letting $r\to+\infty$, we obtain
^$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle=\langle\mathcal{F},\beta\rangle m^{2}+O(m^{2-n})$
as $m\to+\infty$, which matches the formula proven by Stoppa–Szkékelyhidi [SS11, Page 8] for test configurations.
Given an $T$-invariant filtration $\mathcal{F}$ of $(Y,H)$ for a torus $T$ in $\mathrm{Aut}(Y,H)$, we define relative $r$th-Chow weight
(3.4) $\mathrm{Chow}_{r,T}(\mathcal{F})=\mathrm{Chow}_{r}(\mathcal{F})-\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle_{r}}{\langle\beta_{i},\beta_{i}\rangle_{r}}\mathrm{Chow}_{r}(\beta_{i}),$
where $(\beta_{i})$ is any choice of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. In the case of test configurations, Equation (3.4) captures the weight relevant to relative Chow stability.
Lemma 3.7.
We have,
$\mathrm{Chow}_{1,T}(\mathcal{F}_{(r)})\leq\mathrm{Chow}_{r,T}(\mathcal{F}).$
In particular, if $\mathrm{Chow}_{r,T}(\mathcal{F})$ is negative for infinitely many $r$, then $(Y,H)$ is asymptotically relatively Chow unstable with respect to $T$.
Proof.
From [Szé15, Lemma 9], we know the non-relative case
$\mathrm{Chow}_{1}(\mathcal{F}_{(r)})\leq\mathrm{Chow}_{r}(\mathcal{F}).$
This is an equality whenever $\mathcal{F}$ is finitely generated as the approximation $\mathcal{F}_{(r)}$ then is $\mathcal{F}$, viewed as a filtration of $(X,rL)$. Moreover, as ($\mathcal{F}^{\lambda}_{(r)}S_{r}=\mathcal{F}^{\lambda}S_{r}$ for any filtration $\mathcal{F}$, we have $\langle\mathcal{F}_{(r)},\beta_{i}\rangle_{1}=\langle\mathcal{F},\beta_{i}\rangle_{r}$, proving the result. ∎
Assuming now that $p$ is fixed by $T$, we obtain a torus $\widehat{T}$ in $\mathrm{Aut}(\widehat{Y},\widehat{H}_{m})$ which leaves $\widehat{\mathcal{F}}$ invariant. Combining Theorem 3.4 and Proposition 3.5, we obtain the following relative formula:
Proposition 3.8.
We have
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{mr,T}(\mathcal{F})+\Big{(}\nu_{m,T}(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big{)}\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0})$
as $r\to+\infty$ and $m\to+\infty$, where
(3.5) $\nu_{m,T}(p,\mathcal{F})=\nu_{m}(p,\mathcal{F})-\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big{(}\nu_{m}(p,\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big{)},$
for any $\mathbb{Z}$-basis $(\beta_{i})$ of $N_{\mathbb{Z}}(T)$ and $w_{\beta_{i}}(k)=b_{i,0}k^{n+1}+O(k^{n})$.
3.3. Finding destabilizing points.
In order to use the formula in Proposition 3.8 for an instability argument on the blowup, we need to find an invariant point $p$ for which the relative weight in Equation (3.5) is small for large enough $m$. For this, we establish a $T$-invariant version of results of Boucksom [Szé15, Section 8].
Let $S^{\prime}$ be a graded subalgebra of the ring of sections $S=\bigoplus_{k\geq 0}H^{0}(Y,kH)$ not equal to $\mathbb{C}$. Recall the volume of $S^{\prime}$ is the limit
$\mathrm{vol}(S^{\prime})=\lim_{\begin{smallmatrix}k\to\infty\\ S^{\prime}_{k}\neq 0\end{smallmatrix}}\frac{\dim S^{\prime}_{k}}{k^{\kappa}/\kappa!},$
where $\kappa=\mathrm{trdeg}(S^{\prime}/\mathbb{C})-1$ is the Iitaka dimension of $S^{\prime}$.
Definition 3.9.
A graded subalgebra $S^{\prime}$ of $S$ is said to contain an ample series if there exists a decomposition $L=A+E$ of $\mathbb{Q}$-divisors with $A$ ample and $E$ effective such that $H^{0}(X,kA)\subset S_{k}$ for all sufficiently divisible $k$.
The main result of this subsection is the following:
Theorem 3.10.
If $S^{\prime}$ is $T$-invariant, contains an ample series and satisfies $\mathrm{vol}(S^{\prime})<\mathrm{vol}(L)$, then there is a $T$-invariant point $p$ in $X$ and an integer $m_{1}\geq 1$ such that
$S_{mk}^{\prime}\subset H^{0}(X,mkL\otimes\mathcal{I}^{\otimes k}_{p}),$
for all $m\geq m_{1}$ and $k\geq 0$ where $\mathcal{I}_{p}$ is the ideal sheaf of $p$ in $X$.
This is a relative version of [Szé15, Theorem 20] and its proof is based on the following relative versions of [Szé15, Lemma 22] and [Szé15, Lemma 21]:
Lemma 3.11.
Let $S^{\prime}$ and $S^{\prime\prime}$ be two $T$-invariant graded subalgebras of $S$ such that $v(S^{\prime})\geq v(S^{\prime\prime})$ for all $T$-invariant divisorial valuations $v$ of $Y$. Then $\mathrm{vol}(S^{\prime})\leq\mathrm{vol}(S^{\prime\prime})$.
Lemma 3.12.
If there is a $T$-invariant divisorial valuation $v$ on $X$ such that $v(S^{\prime})>0$, then there is a $T$-invariant point $p$ in $X$ such that $\operatorname{ord}_{p}(S^{\prime})>0$.
A divisorial valuation $v$ of $Y$ is a valuation $v:\mathbb{C}(Y)^{*}\to\mathbb{R}$ of the form $v=c\cdot\operatorname{ord}_{E}$ for a real number $c>0$ and a prime divisor $E$ over $Y$, and we define
$v(S^{\prime})=\lim_{k\to+\infty}\frac{\min\{v(f)\;|\;f\in\mathfrak{bs}(S_{k}^{\prime})\setminus\{0\}\}}{k},$
where $\mathfrak{bs}(S_{k}^{\prime})$ is the base-ideal of the linear series $S_{k}^{\prime}$ in $S$. The proofs of Lemmas 3.11 and 3.12 follow verbatim the proofs of [Szé15, Lemma 22] and [Szé15, Lemma 21] noting that the divisorial valuations and their centers constructed in the proofs are invariant, so we omit the details.
Proof of Theorem 3.10.
By Lemma 3.11, there exists a $T$-invariant divisorial valuation $v$ on $X$ such that $v(S^{\prime})>v(S)$. Using that $v(S)=0$, Lemma 3.12 implies that there exists a $T$-invariant point $p$ in $X$ such that $\mathrm{ord}_{p}(S^{\prime})>0$. But this simply means $S_{mk}^{\prime}\subset H^{0}(Y,mkH\otimes\mathcal{I}^{\otimes k}_{p})$ for sufficiently large $m$ and all $k$, as desired. ∎
Remark 3.13.
For test configurations, the existence and construction of a destabilizing point is proven differently in Stoppa–Székelyhidi [SS11], using arguments involving embeddings into projective space.
Let $\mathcal{F}$ be a filtration $(Y,H)$ and write
$e_{\mathrm{max}}(\mathcal{F}):=\lim_{k\to\infty}\frac{\sup\{\lambda\in\mathbb{Z}\;|\;\mathcal{F}^{\lambda}R_{k}\neq 0\}}{k}.$
This value is finite and equals the essential supremum of the concave transform $G_{\mathcal{F}}$ for any choice of Okounkov body [BC11, Section 1]. The desired property that the invariant point in Theorem 3.10 satisfies is the following:
Corollary 3.14.
Let $\Lambda<e_{\mathrm{max}}(\mathcal{F})$ be a real number. There exists a $T$-invariant point $p$ in $X$ such that for any Okounkov body $P$ centered at $p$, there is an integer $m_{1}\geq 1$ such that
$G_{\mathcal{F}}\leq\Lambda\quad\text{on}\quad\Delta_{\frac{1}{m_{1}}}\subset P.$
Proof.
From [BC11, Lemma 1.6] we know that if $\Lambda<e_{\mathrm{max}}(\mathcal{F})$, then the graded subalgebra $S(\Lambda,\mathcal{F})=\bigoplus_{k\geq 0}\mathcal{F}^{\lceil\Lambda k\rceil}S_{k}$ contains an ample series. Moreover, if $\mathcal{F}$ is $T$-invariant, then is $S(\Lambda,\mathcal{F})$ is $T$-invariant by definition. Applying Proposition 3.10 to $S^{\prime}=S(\Lambda,\mathcal{F})$, we obtain the desired upper bound for the concave transform. ∎
3.4. Stability results
In this section, we prove the relevant stability result, Theorem 3.2. Given a $T$-invariant filtration $\mathcal{F}$, we define the relative asymptotic Chow weight
$\mathrm{Chow}_{\infty,T}(\mathcal{F})=\liminf_{r\to+\infty}\mathrm{Chow}_{1,T}(\mathcal{F}_{(r)}),$
where $\mathcal{F}_{(r)}$ is the filtration defined in Example 2.11.
Proposition 3.15.
For any $T$-equivariant good filtration $\mathcal{F}$ of $(X,L)$, we have
$\mathrm{DF}_{T}(\mathcal{F})\geq\mathrm{Chow}_{\infty,T}(\mathcal{F}).$
This is the relative version of [Hat24, Theorem 2.18].
Proof.
From the definition of the relative Donaldson–Futaki invariant, we have $\mathrm{DF}_{T}(\mathcal{F})=\lim_{r\to\infty}\mathrm{Chow}_{r,T}(\mathcal{F})$. Combining this with Lemma 3.7 and the definition of the relative asymptotic Chow weight yields the desired inequality. ∎
We come to the main result of this subsection: for this, assume that the torus $T$ in $\mathrm{Aut}(Y,H)$ is maximal.
Theorem 3.16.
If $(X,L)$ is extremal, then for any $T$-invariant filtration $\mathcal{F}$ of $(X,L)$ with $||\mathcal{F}||_{T}>0$, we have
$\mathrm{Chow}_{\infty,T}(\mathcal{F})>0.$
This is a relative version [Szé15, Theorem 11] allowing extremal metrics and automorphisms. Together with Proposition 3.15, it implies the desired stability result, Theorem 3.2.
Proof.
First, note that it suffices to prove
$\mathrm{Chow}_{\infty}(\mathcal{F}):=\liminf_{r\to+\infty}\mathrm{Chow}_{1}(\mathcal{F}_{(r)})>0,$
for $T$-invariant filtrations $\mathcal{F}$ with $\langle\mathcal{F},N_{\mathbb{Z}}(T)\rangle=0$ and $||\mathcal{F}||>0$. This may be done by replacing $\mathcal{F}$ with the twist $\mathcal{F}_{-\xi(\mathcal{F})}$ for $\xi(\mathcal{F)}$ in $N_{\mathbb{R}}(T)$ as defined in Lemma 2.20. There is a slight subtlety that $\mathcal{F}_{-\xi(\mathcal{F})}$ is an $\mathbb{R}$-filtration, in general; we ignore this and hence assume that $\xi(\mathcal{F})$ is rational, as the argument applies in the same way in general, and $\xi(\mathcal{F})$ is rational in our applications.
Now, assume by contradiction that $\mathrm{Chow}_{\infty}(\mathcal{F})=0$. As $||\mathcal{F}||>0$, the essential infimum $M_{\mathcal{F}}$ of the concave transform $G_{\mathcal{F}}$ is strictly smaller than its average $\overline{G}_{\mathcal{F}}$ over $P$. Note that the values $M_{\mathcal{F}}$ and $\overline{G}_{\mathcal{F}}$ are both independent of choice of Okounkov body. Let
(3.6) $\Lambda_{\mathcal{F}}=\frac{n-2}{n-1}\overline{G}_{\mathcal{F}}+\frac{1}{n-1}M_{\mathcal{F}},$
which will serve as the value in Corollary 3.14 to obtain an upper bound for the concave transform near the origin. Indeed, since $M_{\mathcal{F}}<\overline{G}_{\mathcal{F}}$, we have $\Lambda_{\mathcal{F}}<e_{\mathrm{max}}(\mathcal{F})$ so by Corollary 3.14, there exists a $T$-invariant point $p$ in $X$ such that, after choosing an Okounkov body $P$ centered around $p$, there is an integer $m_{0}\geq 1$, such that
$G_{\mathcal{F}}\leq\Lambda_{\mathcal{F}}\quad\text{on}\quad\Delta_{\frac{1}{m_{0}}}\subset P.$
By assumption, for any $\delta>0$ and $s_{1}\geq 1$, there exists an integer $s\geq s_{1}$, such that $\mathrm{Chow}_{s}(\mathcal{F}_{(s)})<\delta$. We will proceed to set up appropriate choices of $\delta$ and $s_{1}$: let $\delta$ be a real number satisfying
$0<\delta<\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{(n+3)(n-1)}.$
By [Szé15, Lemma 7], there exists an integer $s_{0}\geq 1$ such that for all $s\geq 0$, we have $\overline{G}_{\mathcal{F}}-\delta\leq\overline{G}_{\mathcal{F}_{(s)}}\leq\overline{G}_{\mathcal{F}}$. This gives rise the a priori estimates
$\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}}$ $\leq\frac{nm^{1-n}}{(n-1)!}\Lambda_{\mathcal{F}},$
$\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}}$ $\geq\frac{\sqrt{n}m^{1-n}}{(n-1)!}(M_{\mathcal{F}}-\delta),$
using that the boundary volumes of $\partial_{0}\Delta_{l}$ and $\partial_{0}\Delta_{l}$ are $nl^{n-1}$ and $\sqrt{n}l^{n-1}$, respectively. Moreover, by [Szé15, Lemma 7], the integral formula of the inner product, the boundedness of the weights and orthogonality of $\mathcal{F}$ with respect to $N_{\mathbb{Z}}(T)$, there exists an integer $s_{1}\geq s_{0}$, such that for all $s\geq 0$ and $m\geq m_{0}$, we have
$\bigg{|}\sum_{i=1}^{d}\frac{\langle\mathcal{F}_{(s)},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\mathrm{Chow}_{m}(\beta_{i})\bigg{|}$ $\leq\delta,$
$\bigg{|}\sum_{i=1}^{d}\frac{\langle\mathcal{F}_{(s)},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big{(}\nu_{m}(\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big{)}\bigg{|}$ $\leq\frac{a_{0}\delta}{n-1}.$
With these choices of $\delta$ and $s_{1}$, there exists an integer $s\geq s_{0}$ such that $\mathrm{Chow}_{s}(\mathcal{F}_{(s)})<\delta$. Let $\mathcal{F}_{(s)}^{\prime}$ be the $\overline{G}_{\mathcal{F}_{(s)}}$-shifted filtration of $\mathcal{F}_{(s)}$, given by
$\mathcal{F}_{(s)}^{\prime\lambda}H^{0}(Y,ksH)=\mathcal{F}_{(s)}^{\lambda+\overline{G}_{\mathcal{F}_{(s)}}k}H^{0}(Y,ksH).$
This filtration has the same Chow weights, while the leading coefficient of its weight function vanishes. In particular, the asymptotic expansion of Proposition 3.8 for this filtration takes the form
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}_{(s)}^{\prime}})\leq 2\delta m^{-n}+\nu_{m,T}(p,\mathcal{F}_{(s)}^{\prime})\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0}).$
We now prove that the relative $\nu$-weight in this case is negative: write $\lambda=\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{n-1}>0$, which satisfies both $\Lambda_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-\lambda$ and $M_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-(n-1)\lambda$. Using that $G_{\mathcal{F}_{(s)}^{\prime}}=G_{\mathcal{F}_{(s)}}-\overline{G}_{\mathcal{F}_{(s)}}$, the integrals appearing in the weight satisfy the estimates
$\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}^{\prime}}$ $\leq\frac{nm^{1-n}}{(n-1)!}(\delta-\lambda),$
$\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}^{\prime}}$ $\geq-\frac{\sqrt{n}m^{1-n}}{(n-1)!}(2\delta+(n-1)\lambda).$
Putting this into the $\nu$-weight, we obtain
$\nu_{m,T}(p,\mathcal{F}_{(s)}^{\prime})\leq\frac{a_{0}}{n-1}((n+3)\delta-\lambda),$
noting that the right-hand side is negative as $\delta<\frac{\lambda}{n+3}$ by construction.
Altogether, writing $\eta=\frac{\lambda-(n+3)\delta}{2(n-2)!n-1}>0$, we conclude
(3.7) $\mathrm{Chow}_{r,\widehat{T}}(\widehat{\mathcal{F}_{(s)}^{\prime}})\leq(2\delta-m\eta)m^{-n}+O(m^{1-n}r^{-1}).$
Now, choose $m\geq m_{1}$ to be large enough such $02\delta-m\eta<$ 0 and the blowup $(\widehat{Y},\widehat{sH}_{m})=(\mathrm{Bl}_{p}Y,msH-E)$ is extremal [APS11, Theorem 2.1]. But this is impossible, as [Sey17, Corollary 1.2] implies that the blowup is asymptotically relatively Chow semistable, but Equation (3.7) is negative for infinitely many $r$, contradicting Lemma 3.7. ∎
We used the fact that the torus is maximal in the final step in order to apply [APS11] and [Sey17] (resp. [Mab18, ST21, Has21]).
Remark 3.17.
The result in the case $n=1$ may be reduced to the higher-dimensional result $n=2$ as in [Szé15, Page 18]: any $T$-invariant filtration $\mathcal{F}$ of $(Y,H)$ induces a canonical $T\times T$-invariant filtration $\mathcal{F}^{\prime}$ on the product $(Y\times Y,\mathrm{pr}_{1}^{*}H\otimes\mathrm{pr}_{2}^{*}H)$ satisfying $\mathrm{Chow}_{r,T\times T}(\mathcal{F}^{\prime})=\mathrm{Chow}_{r,T}(\mathcal{F})$.
4. Main results
For the final section, we prove the main results of the introduction, namely Theorem 1.2 and Corollary 1.3. For this, we first examine the relationship between models and filtrations and their twists by one-parameter subgroups.
4.1. Models and filtrations
We start by recalling the construction of a good filtration from two models, due to Hattori [Hat24]. Let $R$ be a discrete valuation ring with fraction field $K$ and residue field $\mathbb{C}$ and let $(X,L)$ be a polarized $K$-variety. Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be models of $(X,L)$ and assume that $L$, $\mathcal{L}$ and $\mathcal{L}^{\prime}$ are genuine line bundles.
[…]
(4.1) $i:H^{0}(\mathcal{Y},k\mathcal{M})\to H^{0}(\mathcal{X},k\mathcal{L}),$
for $\mathcal{M}=\mu^{\prime*}\mathcal{L}^{\prime}-a\mathcal{Y}_{0}$. Choose a uniformizer $t$ in $R$.
Definition 4.1.
The good filtration associated to the models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ is the filtration $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ of $(\mathcal{X}_{0},\mathcal{L}_{0})$, given as
$\mathcal{F}^{\lambda}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})=\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(\mathcal{X},k\mathcal{L})\text{ and }t^{-\lambda}s\in i(H^{0}(\mathcal{Y},k\mathcal{M}))\},$
viewing $t$ as a regular function on $\mathcal{X}$ by pulling back along $\mathcal{X}\to\operatorname{Spec}R$.
In the case where $L$, $\mathcal{L}$ and $\mathcal{L}^{\prime}$ are $\mathbb{Q}$-line bundles, we obtain a filtration of $(\mathcal{X}_{0},r\mathcal{L}_{0})$ for a sufficiently divisible positive integer $r$. The filtration was constructed in [Hat24, Theorem 3.5] in order to prove Theorem 1.1, where it was defined for families over smooth projective curves. However, as the original construction passes through restriction to an affine neighborhood, the original construction coincides with the above definition. Hattori’s good filtration generalizes the filtration constructed by Blum–Xu [BX19] for degenerations of Fano varieties.
The main point of Hattori’s filtration is the following property:
Theorem 4.2.
The filtration $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ is good and satisfies
$w_{\mathcal{F}}(k)=-\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/i(H^{0}(\mathcal{Y},k\mathcal{M}))\big{)},$
for $k\gg 0$. Moreover $||\mathcal{F}||>0$ if and only if the models are not isomorphic.
This was proven in [Hat24, Theorem 3.5 and Corollary 3.9] for families over smooth projective curves. We will see that the same proof strategy applies to models over DVRs.
Proof.
We apply the split short exact sequence of coherent $\mathcal{O}_{\operatorname{Spec}R}$-modules
$0\to t\mathcal{V}/(\mathcal{W}\cap t\mathcal{V})\to\mathcal{V}/\mathcal{W}\to(\mathcal{V}/t\mathcal{V})/\mathrm{im}(\mathcal{W}\to\mathcal{V}/t\mathcal{V})\to 0,$
iteratively to the cases $\mathcal{V}_{\lambda}=t^{\lambda}H^{0}(\mathcal{X},k\mathcal{L})$ and $\mathcal{W}_{\lambda}=t^{\lambda}\mathcal{F}^{-\lambda}H^{0}(\mathcal{X},k\mathcal{L})$ for $\lambda\geq 0$, where $\mathcal{F}^{\lambda}H^{0}(\mathcal{X},k\mathcal{L}):=\{s\;|\;t^{-\lambda}s\in i(H^{0}(\mathcal{Y},k\mathcal{M}))\}$. We have
$t\mathcal{V}_{\lambda}/(\mathcal{W}_{\lambda}\cap t\mathcal{V}_{\lambda})$ $=\mathcal{V}_{\lambda+1}/\mathcal{W}_{\lambda+1},$
$\mathcal{V}_{\lambda}/t\mathcal{V}_{\lambda}$ $=H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}),$
$\mathrm{im}(\mathcal{W}_{\lambda}\to\mathcal{V}_{\lambda}/t\mathcal{V}_{\lambda})$ $=t^{\lambda}\mathcal{F}^{-\lambda}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}),$
so applying the short exact sequence, we obtain
(4.2) $\mathcal{V}_{0}/\mathcal{W}_{0}$ $\simeq\mathcal{V}_{1}/\mathcal{W}_{1}\oplus H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})/t^{0}\mathcal{F}^{-0}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}),$
$\simeq\dots\simeq\bigoplus_{\lambda\geq 0}\big{(}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})/(t^{\lambda}\mathcal{F}^{-\lambda}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}))\big{)}.$
By construction, the left-hand side is $H^{0}(\mathcal{X},kr\mathcal{L})/i(H^{0}(\mathcal{Y},kr\mathcal{M}))$ and the dimension of the right-hand side the weight $w_{\mathcal{F}}(k)$.
For the norm of the filtration, it suffices to show that if the difference $\mu^{*}\mathcal{L}-\mathcal{M}$ supported in $\mathcal{Y}_{0}$ is not $\mathbb{Q}$-linearly trivial, then the concave transform $G_{\mathcal{F}}$ is not constant. Similar to Section 2.1.1, we may assume that $E$ is non-zero, $\mu$-exceptional and $-E$ is $\mu$-ample by taking an ample model.
Assume that $G_{\mathcal{F}}$ is constant and let $f:[0,1]\to\mathbb{R}$ be the function
$f(s)=\mu^{*}\mathcal{L}^{n-1}-\mathcal{M}_{s}^{n+1}$
for $\mathcal{M}_{s}=\mu^{*}\mathcal{L}-sE$, where the intersection number is taken in the sense of, for example, [DR24, Section 2], meaning that $\mu^{*}\mathcal{L}^{n-1}-\mathcal{M}_{s}^{n+1}$ is defined to be the intersection number $(\mu^{*}\mathcal{L}-\mathcal{M}_{s})\cdot(\sum_{j}\mu^{*}\mathcal{L}^{j}\cdot\mathcal{M}^{n-j}_{s})$, where we view $\mu^{*}\mathcal{L}-\mathcal{M}$ is a divisor supported on $\mathcal{Y}_{0}$. This function is smooth with derivative
$f^{\prime}(s)=(n+1)E\cdot\mathcal{M}_{s}^{n},$
which is strictly positive for all $10<s<$ 1 by ampleness. Since the function satisfies $f(0)=0$ and $f(1)=\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1}$, we thus obtain
$\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1}>0.$
On the other hand, the construction of the good filtration yields
$\mathcal{F}^{0}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})=\mathrm{im}(H^{0}(\mathcal{Y},k\mathcal{M})\to H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})).$
The kernel of the map is $H^{0}(\mathcal{Y},k\mathcal{M}-\hat{\mathcal{X}}_{0})$, where $\hat{\mathcal{X}}_{0}$ is the strict transform of $\mathcal{X}_{0}$ along $\mu$. Since $H^{0}(\mathcal{Y},k\mathcal{M}-\hat{\mathcal{X}}_{0})\neq H^{0}(\mathcal{Y},k\mathcal{M})$ for sufficiently large $k$, we have $\mathcal{F}^{0}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})\neq 0$ for sufficiently large $k$, hence there is a rational point $x$ in $P$ such that $G_{\mathcal{F}}(x)=0$.
Moreover, we have $G_{\mathcal{F}}\leq 0$ on the interior of $P$, since the filtration is concentrated in weight $\lambda\leq 0$. By concavity and constancy, this implies that $G_{\mathcal{F}}$ is identically zero almost everywhere.
Combining Equation (2.7) with the first part of the theorem and Equation (2.4), this shows that
(4.3) $\deg\big{(}\lambda_{n+1}(\mathcal{X},\mathcal{L})-\lambda_{n+1}(\mathcal{Y},\mathcal{M})\big{)}=0.$
However, we have $\deg(\lambda_{n+1}(\mathcal{X},\mathcal{L})-\lambda_{n+1}(\mathcal{Y},\mathcal{M}))=\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1}$ using the interpretation of $\lambda_{n+1}$ in terms of Deligne pairings and a result of Boucksom–Eriksson [BE21, Corollary 8.21 and A.23], a contradiction. ∎
Combining Theorem 4.2 with Theorem 2.4, we obtain
(4.4) $\mathrm{DF}(\mathcal{F})=\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})-\mathrm{CM}(\mathcal{X},\mathcal{L}).$
In particular, together with the stability result in [Hat24, Theorem 2.18], one obtains Theorem 1.1 for models defined over DVRs that may not extend to a smooth projective curve.
4.2. Twisting Hattori’s good filtration.
In order to apply Theorem 3.2, we examine twists of Hattori’s good filtration. Let $\beta$ be a $\mathbb{G}_{m,K}$-action on $(X,L)$ and assume that the models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are $\beta$-equivariant. Choose a $\mathbb{G}_{m,R}$-equivariant resolution of indeterminacies $\mathcal{Y}$ of the equivariant birational map $\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$, which makes the embedding in Equation (4.1) $\mathbb{C}^{*}$-equivariant. Write $\beta_{0}$ for the $\mathbb{C}^{*}$-action $(\mathcal{X}_{0},\mathcal{L}_{0})$ obtained by restricting the $\mathbb{G}_{m,R}$-action on $(\mathcal{X},\mathcal{L})$ to the special fiber.
Lemma 4.3.
The filtration $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ is $\beta_{0}$-invariant and satisfies
$w^{2}_{\mathcal{F},\beta_{0}}(k)=-\mathrm{wt}_{\beta}\big{(}H^{0}(\mathcal{X},k\mathcal{L})/i(H^{0}(\mathcal{Y},k\mathcal{M}))\big{)}$
for $k\gg 0$, where $\mathrm{wt}_{\beta}(H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M}))$ denotes the total weight of the $\mathbb{C}^{*}$-action on the quotient $H^{0}(\mathcal{X},k\mathcal{L})/i(H^{0}(\mathcal{Y},k\mathcal{M}))$ induced by $\beta$.
Proof.
Note that $\beta_{0}$-invariance of $\mathcal{F}$ is immediate since $\mathcal{F}^{\lambda}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})$ is the image $\mathbb{C}^{*}$-equivariant map
$\mathcal{F}^{\lambda}H^{0}(\mathcal{X},k\mathcal{L})=\{s\;|\;t^{-\lambda}s\in i(H^{0}(\mathcal{Y},k\mathcal{M}))\}\to H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}).$
Following the definition of the mixed weight-square function, it suffices to prove the identification
$(H^{0}(\mathcal{X},k\mathcal{L})/i(H^{0}(\mathcal{Y},k\mathcal{M})))_{\mu}\simeq\bigoplus_{\lambda\geq 0}t^{\lambda}\big{(}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})_{\mu}/(\mathcal{F}^{-\lambda}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0}))_{\mu}\big{)},$
for large $k$, where the subscripts $\mu$ in $\mathbb{Z}$ indicate weight-space components with respect to the $\mathbb{C}^{*}$-actions. This follows directly from Equation (4.2) as taking weights passes through direct sums and quotients. ∎
As in Section 2.1.2, the choice of uniformizer $t$ in $R$ induces an automorphism $\beta(t):(X,L)\to(X,L)$. The same scheme $\mathcal{Y}$ fits into a new resolution of indeterminacies
${\mathcal{Y}}$${\mathcal{X}^{\prime}}$${\mathcal{X},}$$\scriptstyle{\mu^{\prime}}$$\scriptstyle{\mu_{\beta}}$$\scriptstyle{\psi_{\beta}}$
where $\psi_{\beta}:\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$ is given by $\mathcal{X}^{\prime}_{K}\simeq X\xrightarrow{\beta(t)}X\simeq\mathcal{X}_{K}$ and $\mu^{\prime}$ is as before. The birational morphism $\mu_{\beta}$ still satisfies $\mu^{*}_{\beta}\mathcal{L}=\mu^{\prime*}\mathcal{L}^{\prime}+E-a\mathcal{Y}_{0}$ but gives rise to a twisted embedding
$i_{\beta}:H^{0}(\mathcal{Y},k\mathcal{M})\to H^{0}(\mathcal{X},k\mathcal{L}).$
Lemma 4.4.
We have $i_{\beta}=h_{\beta}\circ i$, where $h_{\beta}:H^{0}(\mathcal{X},k\mathcal{L})\to H^{0}(\mathcal{X},k\mathcal{L})$ is the automorphism defined by
$h_{\beta}(s)=\sum_{\mu\in\mathbb{Z}}t^{-\mu}s_{\mu},$
where we wrote $s=\sum_{\mu\in\mathbb{Z}}s_{\mu}$ with respect to the weight-space decomposition.
Proof.
Over $\operatorname{Spec}K$, we have $\mu_{\beta}=\varphi_{\beta}\circ\mu$, where
$\varphi_{\beta}:\mathcal{X}_{K}\simeq X\xrightarrow{\beta(t)}X\simeq\mathcal{X}_{K}.$
This implies that $i(s)=\varphi^{*}_{\beta}i_{\beta}(s)$ pointwise on $\mathcal{X}_{K}$ for all sections $s$ in $H^{0}(\mathcal{Y},k\mathcal{M})$. The lemma then follows from the fact that the restriction of the automorphism $h_{\beta}$ to $H^{0}(\mathcal{X}_{K},k\mathcal{L}_{K})$ is induced by $\varphi_{\beta}^{-1}$. ∎
Proposition 4.5.
The twist of $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ by $\beta_{0}$ satisfies
$\mathcal{F}^{\lambda}_{\beta_{0}}H^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})=\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(\mathcal{X},k\mathcal{L})\text{ and }t^{-\lambda}s\in i_{\beta}(H^{0}(\mathcal{Y},k\mathcal{M}))\}.$
In particular, the twisted filtration $\mathcal{F}_{\beta_{0}}$ coincides with the good filtration associated to the twisted model $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$.
Proof.
Writing subscripts for the weight-space decomposition components, a direct computation shows that
$\mathcal{F}^{\lambda}_{\beta_{0}}H^{0}(\mathcal{X},kr\mathcal{L})$ $=\bigoplus_{\mu\in\mathbb{Z}}\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})\text{ and }t^{\mu-\lambda}s\in i(H^{0}(kr\mathcal{M}))\}_{\mu},$
$=\bigoplus_{\mu\in\mathbb{Z}}\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})_{\mu}\text{ and }t^{\mu-\lambda}t^{-\mu}s\in i_{\beta}(H^{0}(kr\mathcal{M}))\},$
$=\bigoplus_{\mu\in\mathbb{Z}}\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})\text{ and }t^{-\lambda}s\in i_{\beta}(H^{0}(kr\mathcal{M}))\}_{\mu},$
=$=\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})\text{ and }t^{-\lambda}s\in i_{\beta}(H^{0}(kr\mathcal{M}))\},$
as desired. ∎
Lastly, we characterize positivity of the reduced norm of a good filtration induced by two equivariant models. Let $T$ be a torus in $\mathrm{Aut}_{K}(X,L)$. By Lemma 4.3, if $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are $T$-equivariant, then the good filtration $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ is $T_{0}$-invariant, where $T_{0}$ in $\mathrm{Aut}(\mathcal{X}_{0},\mathcal{L}_{0})$ is the restriction of the torus $T_{R}$ in $\mathrm{Aut}_{R}(\mathcal{X},\mathcal{L})$ to the special fiber.
Proposition 4.6.
We have $||\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ one-parameter subgroups $\beta$ of $T$.
Proof.
From Lemma 2.20, we know that the infimum over $N_{\mathbb{R}}(T)$ in the definition of the reduced norm is attained by $\xi(\mathcal{F})=\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$ where $(\beta_{i})$ is any $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. As $\mathcal{F}$ is good, the inner product $\langle\mathcal{F},\beta_{i}\rangle$ is a rational number, so $\xi(\mathcal{F})$ lies in $N_{\mathbb{Q}}(T)$. We obtain
$||\mathcal{F}||_{T_{0}}=\min_{\xi\in N_{\mathbb{Q}}(T)}||\mathcal{F}_{\xi}||,$
where the twist $\mathcal{F}_{\xi}$ may be viewed as a $\mathbb{Z}$-filtration of $(\mathcal{X}_{0},r\mathcal{L}_{0})$ for a sufficiently divisible positive integer $r$. Applying Theorem 4.2, we obtain $||\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\xi},\mathcal{L}_{\xi})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all rational twists $\xi$ in $N_{\mathbb{Q}}(T)$. But this is equivalent to $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ in $N_{\mathbb{Z}}(T)$ as we may simply scale a given automorphism. ∎
4.3. Proof of the main results
We start by restating the first main result, Theorem 1.2. Let $T$ be a maximal torus in $\mathrm{Aut}_{K}(X,L)$.
Theorem 4.7.
Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be $T$-equivariant models of $(X,L)$ such that the restricted torus $T_{0}$ in $\mathrm{Aut}(\mathcal{X}_{0},\mathcal{L}_{0})$ is maximal. If $(\mathcal{X}_{0},\mathcal{L}_{0})$ smooth and extremal, then there exists a rational one-parameter subgroup $\xi$ of $T$, such that
$\mathrm{CM}(\mathcal{X}_{\xi},\mathcal{L}_{\xi})\leq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}),$
where the inequality is strict if and only if $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all one-parameter subgroups $\beta$ of $T$.
Proof.
By invariance of the difference of CM degrees and the Donaldson–Futaki invariant under scaling, we may assume that $L$, $\mathcal{L}$ and $\mathcal{L}$ are all line bundles. The good filtration $\mathcal{F}=\mathcal{F}_{(\mathcal{X},\mathcal{L}),(\mathcal{X}^{\prime},\mathcal{L}^{\prime})}$ is $T_{0}$-invariant and, following Proposition 4.5 and Equation (4.4), we have
$\mathrm{DF}_{T_{0}}(\mathcal{F})=\mathrm{DF}(\mathcal{F}_{\xi})=\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})-\mathrm{CM}(\mathcal{X}_{\xi},\mathcal{L}_{\xi}),$
for $\xi=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$, where $(\beta_{i})$ is any of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. As $(\mathcal{X}_{0},\mathcal{L}_{0})$ is extremal, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$ with strict inequality if and only if $||\mathcal{F}||_{T_{0}}>0$ by Theorem 3.2. The result follows by Proposition 4.6. ∎
Remark 4.8.
The proof reveals that we may explicitly give the rational one-parameter subgroup $\xi$ as $-\xi(\mathcal{F})=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$ from Lemma 2.20. Geometrically, it may be viewed as the orthogonal projection of the filtration $\mathcal{F}$ to $N_{\mathbb{Q}}(T)$ whose coefficients may be expressed geometrically in terms of Hilbert- and weight polynomials of the models, following Theorem (4.2) and Lemma 4.3.
By replacing $\mathrm{DF}_{T_{0}}(\mathcal{F})$ by $\mathrm{DF}(\mathcal{F})$ and applying Corollary (3.3) with Theorem 4.2 (which uses the vanishing of the Futaki invariant for cscK manifolds [Fut83]), we obtain Corollary 1.3 for cscK metrics.